OpenStudy (anonymous):

Find the surface area of a cylinder with d = 12 in: h = 14 in

OpenStudy (lucaz):

base is a cricle, area of circle is pi*r^2

OpenStudy (lucaz):

radius is d/2 = 6 base = pi*6^2 in

OpenStudy (lucaz):

sorry, you have to add the basse to the 'lateral' area

OpenStudy (lucaz):

so it's 2*base + circunference*height

OpenStudy (lucaz):

circunference is pi*d C=pi*12in SA = 2(pi*6^2)+(pi*12) = 84pi

OpenStudy (anonymous):

113.04 sq. in. B. 640.56 sq. in. C. 527.52 sq. in. D. 753.6 sq. in. these are my choices

OpenStudy (lucaz):

it's D, let me do it again..